If escape velocity on a planet is 11.2 km per second then on planet which has mass 4 times that of earth and acceleration due to gravity equal to earth the escape velocity will be?

Grade:12th pass

1 Answers

Norman Manchu
6 years ago
The escape velocity of earth is about 11km|s thus for earth-
11.2=\sqrt\frac{2GM}{R}----(1)
For planet-
V=\sqrt\frac{2G(4M)}{R}----(2)
Divide equations hence you'll get v=22.4km|s